# Online Housing Market

### v1

@@ 5 Dynamic Serial Dictatorship Procedure
In this section, I explore a dynamic version of Algorithm 1, where the choice made by the serial dictatorship procedure can be revised when new items arrive to the market. The final decision about the item allocated to an agent is made when this agent leaves the market. As described in Algorithm 2, a serial dictatorship procedure is used to decide which item will be allocated to the agent leaving the market. Each agent appearing before the agent leaving the market in the sequence selects an item among the remaining ones, and after that, the agent leaving the market picks the best remaining item left by the others. Note that, except for the agent that leaves the market, no item is allocated to an agent. We can see the choices made by these agents remaining in the market as reservations that save all selected items to be picked by the leaving agent. However, the remaining agents may leave the market with other items, including those that will appear in the market later on. Note that Algorithm 2 makes use of the ascending departure permutation $\delta$ presented in the previous section, in order to treat the agents by increasing departure times, which is required for the algorithm to be online.
@@
7:     $T\leftarrow\emptyset$Set of reserved items.}
@@
As with Algorithm 1, Algorithm 2 should use an online permutation function as input to be online. However, contrary to Algorithm 1, Algorithm 2 does not require the use of $SOC$ permutation functions to run properly or to be online. However, incentive compatibility is more difficult to obtain since the procedure may ask multiple times an agent about her most favorite item among the remaining ones, and agents may have greater incentive to misreport. Therefore, Algorithm 2 requires the use of specific permutation functions to incentivize agents to truthfully report their preferences. Here is a definition describing this type of permutation functions.
@@
Permutation function $\Pi$ is dynamic online compatible ($DOC$), if for any three agents $i,j$ and $k$ such that $a_{i}<d_{k}<d_{i}$, $d_{k}<d_{j}$, $d_{i}>a_{j}$, agent $k$ is ranked behind agent $i$ in $\Pi(I_{<d_{k}})$, and either $a_{j}>d_{k}$ or $j$ is also ranked behind agent $i$ in $\Pi(I_{<d_{k}})$, then agent $j$ is ranked behind agent $i$ in $\Pi(I_{<\min{\{d_{i},d_{j}\}}})$.
To put it in simple terms, the order over the agents that are ranked before the leaving agent should not change afterward. This means that an agent $i$ that is ranked in $\Pi(I_{<d_{k}})$ before the leaving agent $k$, should not be ranked later on after another agent $j$ that was either ranked after $i$ in $\Pi(I_{<d_{k}})$ or not even present before agent $k$ leaves the market. The intuition behind this formal definition is that the order in which agents reserve their items during Algorithm 2 should not change from iteration to iteration, because otherwise they may adopt a strategic behavior. For example, when agent $k$ is leaving the market, agent $i$ may strategically reserve an item for agent $j$ in order to avoid that later on when agent $j$ chooses before agent $i$, she picks or reserves an item that is better for agent $i$ [^4] The full explanation is provided in the proof of Proposition 9. The following proposition characterizes the $DOC$ permutation functions as the ones which make Algorithm 2 $WIC$.
@@
The exchange algorithm described in Algorithm 2 is $WIC$ if and only if the permutation function $\Pi$ is $DOC$.
@@
Assume first that $\Pi$ is not $DOC$. This means that there is an instance $I$ and three agents $i$, $j$, and $k$ such that $k$ leaves earlier than $i$ and $j$, $\Pi(I_{<d_{k}})$ orders $i$ before $j$ and $k$, whereas either $\Pi(I_{<d_{j}})$ or $\Pi(I_{<d_{i}})$ orders $j$ before $i$, depending on who between $j$ and $i$ leaves first. Without loss of generality, I can assume that the preferences of the agents for this instance are as follows. [^5] Note that I assumed earlier in the paper that the outcome of a permutation function is independent of the preferences of the agents. The preferences of agent $i$ are $e_{k}\succ_{i}e_{i}\succ_{i}e_{j}\succ_{i}\ldots$, where all the other items are ranked arbitrarily after these three items. The preferences of agent $j$ are $e_{k}\succ_{j}e_{j}\succ_{j}e_{i}\succ_{j}\ldots$, and the preferences of agent $k$ are $e_{i}\succ_{k}e_{k}\succ_{k}e_{j}\succ_{k}\ldots$. Finally, the preferences $\succ_{\ell}$ of any other agent $\ell$ are such that $e_{\ell}$ appears in first position, whereas the other items appear in an arbitrary order.
@@
If I apply Algorithm 2 to this instance $I$, then each other agent $\ell$ other than $i$, $j$, and $k$ will always choose her own item. Furthermore, when agent $k$ leaves the market, agent $i$, who chooses before agents $k$ and $j$ (if $j$ has already arrived at $d_{k}$), reserves $e_{k}$. Independently of the presence of agent $j$ in $I_{<d_{k}}$, or her choice if she is ranked before agent $k$ in $\Pi(I_{<d_{k}})$, agent $k$ selects $e_{i}$, and $e_{i}$ is assigned to agent $k$ when she leaves the market. Then, in both cases where either agent $i$ or agent $j$ leaves first, agent $j$ selects before agent $i$ and chooses $e_{k}$. When agent $i$ leaves the market, she receives $e_{j}$, her third choice.
@@
Now, assume that agent $i$ reveals other preferences, say $\succ^{\prime}_{i}$, instead of her true preferences. These new preferences are such that $e_{i}\succ^{\prime}_{i}e_{k}\succ^{\prime}_{i}e_{j}\succ^{\prime}_{i}\ldots$. Let $I^{\prime}$ denote the copy of $I$ where agent $i$ reveals $\succ^{\prime}_{i}$ instead of $\succ_{i}$, and the other agents keep the same preferences. If I apply Algorithm 2 to $I^{\prime}$, then agent $i$ reserves $e_{i}$ instead of $e_{k}$ when agent $k$ leaves the market, and agent $k$ picks $e_{k}$ or $e_{j}$, depending on the position of $j$ in $\Pi(I_{<d_{k}})$. Then, at time $\min\{d_{i},d_{j}\}$, when either $i$ or $j$ leaves the market, agent $j$ selects either $e_{k}$ or $e_{j}$, depending on the item assigned earlier to agent $k$, and agent $i$ receives ultimately $e_{i}$, which is strictly better than $e_{j}$ according to her true preferences $\succ_{i}$. Therefore, agent $i$ has an incentive to misreport her preferences.
@@
Now, I am going to show that if $\Pi$ is $DOC$, then Algorithm 2 is $WIC$. To do so, I first show that for any agent $i$, the set of items selected by the agent ranked before $i$ in $\Pi$ will not change if agent $i$ changes her revealed preferences. Let’s consider the first iteration of the “for” loop of Algorithm 2 where agent $i$ has to select an item. Note that this iteration will occur at the same time whatever preferences are revealed by agent $i$. If this is the iteration corresponding to $d_{i}$, where agent $i$ leaves the market, then she has no incentive to misreport her preferences. Therefore, I can focus on the cases where the iteration does not correspond to $d_{i}$, and agent $i$ remains in the market in the next iterations. Let $K$ denote the set of items remaining in the market and not selected by one of the agents ranked before agent $i$ in $\Pi$. I am going to show that in the next iteration, no agent ranked before $i$ in $\Pi$ will select an item of $K$.
@@
Let me show this by induction on the relative rank of the agent among the ones reserving before $i$. I start with the first agent, who selects her most favorite item among the remaining ones. During the previous iteration, she reserved an item which is not part of $K$. So this item must still be available. During this new iteration, she may prefer another item which was not present during the previous iteration, and leave unreserved the item that she chose before, but she will not select an item from $K$.
@@
Now, let me consider that the first $k-1$ ranked before agent $i$ did not select an item from $K$, and let me show that the agent at rank $k$ also does not select an item from $K$. Note that each of the $k-1$ first agents reserved during this iteration either the item that they reserved during the previous iteration, or an item that arrived later than the previous iteration, or an item reserved during the previous iteration by an agent ranked at a better position but who chose another item during this iteration and left the previously reserved item free. This means that the item reserved during the previous iteration by the agent at position $k$ is still available, and she can reserve it again, or reserve any other item that was not available during her turn in the previous iteration that she prefers. Thus, she will not select an item from $K$.
@@
This completes the induction and shows that no agent ranked before $i$ in $\Pi$ will be interested in selecting an item from $K$ in the next iteration of the “for” loop of Algorithm 2. By applying the same reasoning, and by considering for each new iteration a larger set $K$ of items not chosen by the agents ranked before $i$, I can show that from iteration to iteration, no agent ranked before $i$ will select an item accessible to $i$ when it is her turn to reserve. This implies that the choice made by $i$ on the item that she reserves during each iteration has no consequence on the set of items selected by the agents who are ranked before her in $\Pi$. Therefore, she has no incentive to misreport her true preferences in order to reserve another item instead of her most favorite one, because if she does so, another agent ranked after her may pick it and leave the market with it. ∎
It is easy to check that $SOC$ is stronger than $DOC$.
@@
Any permutation function that is $SOC$ is also $DOC$.
@@
Let me first show that Algorithm 2 using the ascending arrival permutation $\alpha$ as input is $d$-$IC$. Note first that the ascending arrival permutation $\alpha$ is $SOC$, and therefore by Proposition 9 and Observation 10, I know that no agent has an incentive to misreport her preferences. So I only need to show that no agent has an incentive to misreport her departure time. Note also that as soon as the arrival does not change, changing the departure time of an agent has no impact on the ranking dictated by the ascending arrival permutation $\alpha$. I am going to show that if the departure time of an agent, say $i$, is delayed after a new event, which may be the departure of an agent or the arrival of a new agent, then her most favorite available item is either the same or even better.
@@
Let me now run the dynamic serial dictatorship procedure with the ascending arrival permutation $\alpha$. At $d_{2}$, agent 1 is the first to choose, and she reserves $e_{3}$. Then, agent 2 is the second to choose, and she picks $e_{1}$. Item $e_{1}$ is therefore allocated to agent 3, and all other items remain unallocated. At $d_{1}$, agent 1 is the first to choose once again and she picks $e_{4}$ this time, which was not available at $d_{2}$. Therefore, item $e_{4}$ is allocated to agent 1. At $d_{3}$, agent 3 is the first to choose, and she picks item $e_{2}$. Finally, at $d_{4}$, agent 4 picks the only remaining item, $e_{3}$. The resulting allocation is therefore $M^{\prime\prime}$, as displayed in Figure 2. Note that $M^{\prime\prime}$ Pareto-dominates $M^{\prime}$, which is the result of the static serial dictatorship procedure using the same ascending arrival permutation $\alpha$. However, this will not necessarily be the case each time, even though the result of the dynamic version of serial dictatorship will never be Pareto-dominated by the result of the static one.
@@
Let me consider instance $I_{<d_{1}}$ where agent 3 is removed. Once again, because the algorithm is $a$-$IC$, agent 2 should receive item $e_{2}$, since otherwise she can delay her arrival time after the departure of agent 1 and keep her item. This implies that item $e_{1}$ is assigned to agent 1 in $I_{<d_{1}}$. Note that this implies that $\Pi_{1}(I_{<d_{1}})=2$ and $\Pi_{2}(I_{<d_{1}})=1$ hold since $e_{2}$ is preferred to $e_{1}$ by agent 1 and would be chosen if it is not reserved by agent 2. Furthermore, since $e_{3}$ should be assigned to agent 3 in $I$ and $e_{3}$ is the most favorite item of agent 2, agent 3 should be able to reserve it, and therefore be ranked before agent 2 in $\Pi(I_{<d_{2}})=\Pi(I)$. But the fact that $\Pi$ is $a$-$IC$ implies that it is also $WIC$, and by Proposition 9 I know that $\Pi$ should be $DOC$. This leads to a contradiction since in the instance described earlier, we had agent 2 ranked before agent 1 in $\Pi_{I_{<d_{1}}}$, whereas agent 2 is ranked after an agent arriving later than $d_{1}$, namely agent 3, in $\Pi(I_{<d_{2}})$.

### v2

@@ 5 Dynamic Serial Dictatorship Procedure
In this section, we explore a dynamic version of Algorithm 1, where the choice made by the serial dictatorship procedure can be revised when new items arrive to the market. The final decision about the item allocated to an agent is made when this agent leaves the market. As described in Algorithm 2, a serial dictatorship procedure is used to decide which item will be allocated to the agent leaving the market. Each agent appearing before the agent leaving the market in the sequence selects an item among the remaining ones, and after that, the agent leaving the market picks the best remaining item left by the others. Note that, except for the agent that leaves the market, no item is allocated to an agent. We can see the choices made by these agents remaining in the market as reservations that save all selected items to be picked by the leaving agent. However, the remaining agents may leave the market with other items, including those that will appear in the market later on. Note that Algorithm 2 makes use of the ascending departure permutation $\delta$ presented in the previous section, in order to treat the agents by increasing departure times, which is required for the algorithm to be online.
@@
7:     $T\leftarrow\emptyset$Set of items reserved in this phase.}
@@
As with Algorithm 1, Algorithm 2 should use an online permutation function as input to be online. However, contrary to Algorithm 1, Algorithm 2 does not require the use of a $PS$ permutation function to run properly or to be online. However, incentive compatibility is more difficult to obtain since the procedure may ask multiple times an agent about her most favorite item among the remaining ones, and agents may have greater incentive to misreport. Therefore, Algorithm 2 requires the use of specific permutation functions to incentivize agents to truthfully report their preferences. Here is a definition describing this type of permutation functions.
@@
Permutation function $\Pi$ is prefix-stable extended ($PFE$), if for any three agents $i,j$ and $k$ such that $a_{i}<d_{k}<d_{i}$, $d_{k}<d_{j}$, $d_{i}>a_{j}$, agent $k$ is ranked behind agent $i$ in $\Pi(I_{<d_{k}})$, and either $a_{j}>d_{k}$ or $j$ is also ranked behind agent $i$ in $\Pi(I_{<d_{k}})$, then agent $j$ is ranked behind agent $i$ in $\Pi(I_{<\min{\{d_{i},d_{j}\}}})$.
To put it in simple terms, the order over the agents that are ranked before the leaving agent should not change afterward. [^8] As the name suggests, the prefix-stable extended concept is closely related to the notion of prefix stability. This means that an agent $i$ that is ranked in $\Pi(I_{<d_{k}})$ before the leaving agent $k$, should not be ranked later on after another agent $j$ that was either ranked after $i$ in $\Pi(I_{<d_{k}})$ or not even present before agent $k$ leaves the market. The intuition behind this formal definition is that the order in which agents reserve their items during Algorithm 2 should not change from iteration to iteration, because otherwise they may adopt a strategic behavior. For example, when agent $k$ is leaving the market, agent $i$ may strategically reserve an item for agent $j$ in order to avoid that later on when agent $j$ chooses before agent $i$, she picks or reserves an item that is better for agent $i$ [^9] The full explanation is provided in the proof of Proposition 9. The following proposition characterizes the $PFE$ permutation functions as the ones which make Algorithm 2 $WIC$.
@@
The exchange algorithm described in Algorithm 2 is $WIC$ if and only if the permutation function $\Pi$ is $PFE$.
@@
Assume first that $\Pi$ is not $PFE$. This means that there is an instance $I$ and three agents $i$, $j$, and $k$ such that $k$ leaves earlier than $i$ and $j$, $\Pi(I_{<d_{k}})$ orders $i$ before $j$ and $k$, whereas either $\Pi(I_{<d_{j}})$ or $\Pi(I_{<d_{i}})$ orders $j$ before $i$, depending on who between $j$ and $i$ leaves first. Without loss of generality, we can assume that the preferences of the agents for this instance are as follows. [^10] Note that we assumed earlier in the paper that the outcome of a permutation function is independent of the preferences of the agents. The preferences of agent $i$ are $e_{k}\succ_{i}e_{i}\succ_{i}e_{j}\succ_{i}\ldots$, where all the other items are ranked arbitrarily after these three items. The preferences of agent $j$ are $e_{k}\succ_{j}e_{j}\succ_{j}e_{i}\succ_{j}\ldots$, and the preferences of agent $k$ are $e_{i}\succ_{k}e_{k}\succ_{k}e_{j}\succ_{k}\ldots$. Finally, the preferences $\succ_{\ell}$ of any other agent $\ell$ are such that $e_{\ell}$ appears in first position, whereas the other items appear in an arbitrary order.
@@
If we apply Algorithm 2 to this instance $I$, then each other agent $\ell$ other than $i$, $j$, and $k$ will always choose her own item. Furthermore, when agent $k$ leaves the market, agent $i$, who chooses before agents $k$ and $j$ (if $j$ has already arrived at $d_{k}$), reserves $e_{k}$. Independently of the presence of agent $j$ in $I_{<d_{k}}$, or her choice if she is ranked before agent $k$ in $\Pi(I_{<d_{k}})$, agent $k$ selects $e_{i}$, and $e_{i}$ is assigned to agent $k$ when she leaves the market. Then, in both cases where either agent $i$ or agent $j$ leaves first, agent $j$ selects before agent $i$ and chooses $e_{k}$. When agent $i$ leaves the market, she receives $e_{j}$, her third choice.
@@
Now, assume that agent $i$ reveals other preferences, say $\succ^{\prime}_{i}$, instead of her true preferences. These new preferences are such that $e_{i}\succ^{\prime}_{i}e_{k}\succ^{\prime}_{i}e_{j}\succ^{\prime}_{i}\ldots$. Let $I^{\prime}$ denote the copy of $I$ where agent $i$ reveals $\succ^{\prime}_{i}$ instead of $\succ_{i}$, and the other agents keep the same preferences. If we apply Algorithm 2 to $I^{\prime}$, then agent $i$ reserves $e_{i}$ instead of $e_{k}$ when agent $k$ leaves the market, and agent $k$ picks $e_{k}$ or $e_{j}$, depending on the position of $j$ in $\Pi(I_{<d_{k}})$. Then, at time $\min\{d_{i},d_{j}\}$, when either $i$ or $j$ leaves the market, agent $j$ selects either $e_{k}$ or $e_{j}$, depending on the item assigned earlier to agent $k$, and agent $i$ receives ultimately $e_{i}$, which is strictly better than $e_{j}$ according to her true preferences $\succ_{i}$. Therefore, agent $i$ has an incentive to misreport her preferences.
@@
Now, we are going to show that if $\Pi$ is $PFE$, then Algorithm 2 is $WIC$. To do so, we first show that for any agent $i$, the set of items selected by the agent ranked before $i$ in $\Pi$ will not change if agent $i$ changes her revealed preferences. Let’s consider the first iteration of the “for” loop of Algorithm 2 where agent $i$ has to select an item. Note that this iteration will occur at the same time whatever preferences are revealed by agent $i$. If this is the iteration corresponding to $d_{i}$, where agent $i$ leaves the market, then she has no incentive to misreport her preferences. Therefore, we can focus on the cases where the iteration does not correspond to $d_{i}$, and agent $i$ remains in the market in the next iterations. Let $K$ denote the set of items remaining in the market and not selected by one of the agents ranked before agent $i$ in $\Pi$. We are going to show that in the next iteration, no agent ranked before $i$ in $\Pi$ will select an item of $K$.
@@
Let us show this by induction on the relative rank of the agent among the ones reserving before $i$. We start with the first agent, who selects her most favorite item among the remaining ones. During the previous iteration, she reserved an item which is not part of $K$. So this item must still be available. During this new iteration, she may prefer another item which was not present during the previous iteration, and leave unreserved the item that she chose before, but she will not select an item from $K$.
@@
Now, let us consider that the first $k-1$ ranked before agent $i$ did not select an item from $K$, and let us show that the agent at rank $k$ also does not select an item from $K$. Note that each of the $k-1$ first agents reserved during this iteration either the item that they reserved during the previous iteration, or an item that arrived later than the previous iteration, or an item reserved during the previous iteration by an agent ranked at a better position but who chose another item during this iteration and left the previously reserved item free. This means that the item reserved during the previous iteration by the agent at position $k$ is still available, and she can reserve it again, or reserve any other item that was not available during her turn in the previous iteration that she prefers. Thus, she will not select an item from $K$.
@@
This completes the induction and shows that no agent ranked before $i$ in $\Pi$ will be interested in selecting an item from $K$ in the next iteration of the “for” loop of Algorithm 2. By applying the same reasoning, and by considering for each new iteration a larger set $K$ of items not chosen by the agents ranked before $i$, we can show that from iteration to iteration, no agent ranked before $i$ will select an item accessible to $i$ when it is her turn to reserve. This implies that the choice made by $i$ on the item that she reserves during each iteration has no consequence on the set of items selected by the agents who are ranked before her in $\Pi$. Therefore, she has no incentive to misreport her true preferences in order to reserve another item instead of her most favorite one, because if she does so, another agent ranked after her may pick it and leave the market with it. ∎
Note that, in essence, the definition of $PFE$ is quite similar to Definition 5 of prefix-stability. However, to establish the equivalence in Proposition 9, we had to extend the definition to permutations where the positions of agents who have already left the market are not fixed (since they are ignored by Algorithm 2). It is therefore easy to verify that the notion of $PS$ is slightly stronger than that of $PFE$.
@@
Any permutation function that is $PS$ is also $PFE$.
@@
Let us first show that Algorithm 2 using the ascending arrival permutation $\alpha$ as input is $d$-$IC$. Note first that the ascending arrival permutation $\alpha$ is $PS$, and therefore by Proposition 9 and Observation 10, we know that no agent has an incentive to misreport her preferences. So we only need to show that no agent has an incentive to misreport her departure time. Note also that as soon as the arrival does not change, changing the departure time of an agent has no impact on the ranking dictated by the ascending arrival permutation $\alpha$. We are going to show that if the departure time of an agent, say $i$, is delayed after a new event, which may be the departure of an agent or the arrival of a new agent, then her most favorite available item is either the same or even better.
@@
Let us now run the dynamic serial dictatorship procedure with the ascending arrival permutation $\alpha$. At $d_{2}$, agent 1 is the first to choose, and she reserves $e_{3}$. Then, agent 2 is the second to choose, and she picks $e_{1}$. Item $e_{1}$ is therefore allocated to agent 2, and all other items remain unallocated. At $d_{1}$, agent 1 is the first to choose once again and she picks $e_{4}$ this time, which was not available at $d_{2}$. Therefore, item $e_{4}$ is allocated to agent 1. At $d_{3}$, agent 3 is the first to choose, and she picks item $e_{2}$. Finally, at $d_{4}$, agent 4 picks the only remaining item, $e_{3}$. The resulting allocation is therefore $M^{\prime\prime}$, as displayed in Figure 2. Note that $M^{\prime\prime}$ Pareto-dominates $M^{\prime}$, which is the result of the static serial dictatorship procedure using the same ascending arrival permutation $\alpha$. However, this will not necessarily be the case each time, even though the result of the dynamic version of serial dictatorship will never be Pareto-dominated by the result of the static one.
@@
Let us consider instance $I_{<d_{1}}$ where agent 3 is removed. Once again, because the algorithm is $a$-$IC$, agent 2 should receive item $e_{2}$, since otherwise she can delay her arrival time after the departure of agent 1 and keep her item. This implies that item $e_{1}$ is assigned to agent 1 in $I_{<d_{1}}$. Note that this implies that $\Pi_{1}(I_{<d_{1}})=2$ and $\Pi_{2}(I_{<d_{1}})=1$ hold since $e_{2}$ is preferred to $e_{1}$ by agent 1 and would be chosen if it is not reserved by agent 2. Furthermore, since $e_{3}$ should be assigned to agent 3 in $I$ and $e_{3}$ is the most favorite item of agent 2, agent 3 should be able to reserve it, and therefore be ranked before agent 2 in $\Pi(I_{<d_{2}})=\Pi(I)$. But the fact that $\Pi$ is $a$-$IC$ implies that it is also $WIC$, and by Proposition 9 we know that $\Pi$ should be $PFE$. This leads to a contradiction since in the instance described earlier, we had agent 2 ranked before agent 1 in $\Pi_{I_{<d_{1}}}$, whereas agent 2 is ranked after an agent arriving later than $d_{1}$, namely agent 3, in $\Pi(I_{<d_{2}})$.